Use the present value tables to calculate the issue price of a $500,000 bond issue in each of the following independent cases. Assume that the bond was issued on January 1, 2010 and that interest is paid quarterly on March 31, June 30 September 30 and December 31.

A)A 10-year, 8 percent bond issue; the market interest rate is 12 percent

B)A 10-year, 12 percent bond issue; the market interest rate is 8 percent

C)A 5-year, 12 percent bond issue; the market interest rate is 8 percent

D)A 5-year, 8 percent bond issue; the market interest rate is 12 percent

3. Using the information from question # 1a and 1b above, prepare the amortization schedules and make the entry to record the interest expense for the years 2010, 2011 & 2012

4. Using the information from question #3, make the journal entry to record the retirement on bonds on January 1, 2013 at the following independent prices
A) 95
B) 106
Answer #1

Based on the information available in the question, we can calculate the Issue price of the bond as follows:-

A.)

Table values based on
n= 40
i= 3%
Cash flow Table Value Amount Present Value
Par(Maturity value) 500,000 0.3644      182,200
Interest(Annuity)               10,000 23.1148      231,148
Price of Bonds      413,348

I = 12% / 4 times = 3%

N = 10 years * 4 payments per year = 40 payments

Interest(Annuity)= $500,000 * 2%(Stated quarterly rate) = $10,000

B.)

Table values based on
n= 40
i= 2%
Cash flow Table Value Amount Present Value
Par(Maturity value)            500,000 0.5102      255,100
Interest(Annuity)               15,000 27.3555      410,333
Price of Bonds      665,433

I = 8% / 4 times = 2%

N = 10 years * 4 payments per year = 40 payments

Interest (Annuity) = $500,000 * 3%(Stated quarterly rate) = $15,000

C.)

Table values based on
n= 20
i= 2%
Cash flow Table Value Amount Present Value
Par(Maturity value)            500,000 0.6944      347,200
Interest(Annuity)               15,000 16.3514      245,271
Price of Bonds      592,471

I = 8% / 4 times = 2%

N = 5 years * 4 payments per year = 20 payments

Interest (Annuity) = $500,000 * 3%(Stated quarterly rate) = $15,000

D.)

Table values based on
n= 20
i= 3%
Cash flow Table Value Amount Present Value
Par(Maturity value)            500,000 0.5787      289,350
Interest(Annuity)               10,000 14.8775      148,775
Price of Bonds      438,125

  I = 12% / 4 times = 3%

N = 5 years * 4 payments per year = 20 payments

Interest (Annuity) = $500,000 * 2%(Stated quarterly rate) = $10,000

Please be informed that we have answered the first 4 questions HOMEWORKLIB answering guidelines. Kindly post the remaining questions separately so that we can answer them as well. All the best and please let me know if you have any questions via comments :)